(x-4y-9)dx+(4x+y-2)dy=0
1
Expert's answer
2020-06-15T19:14:25-0400

"(x-4y-9)dx+(4x+y-2)dy=0"

Find the solution of system equation

"\left\{\begin{matrix}
 x-4y=9 \\
 4x+y=2
\end{matrix}\right.\\
\left\{\begin{matrix}
 x=1 \\
 y=-2
\end{matrix}\right."

Make a replacement

"\left\{\begin{matrix}
 x=t+1 \\
 y=z(t)-2
\end{matrix}\right."

and input in equation

"z'=-\frac{t+1-4(z-2)-9}{4(t+1)+z-2-2}\\
z'=-\frac{t-4z}{4t+z}"

Make a replacement

"z=t\cdot u(t)\\
tu'+u=-\frac{t-4tu}{4t+tu}\\
tu'=\frac{4u-1}{u+4}-u\\
tu'=\frac{-u^2-1}{u+4}\\
\frac{u+4}{u^2+1}du=-\frac{dt}{t}\\
\int{\frac{u}{u^2+1}+\frac{4}{u^2+1}}du=-\int\frac{dt}{t}\\
\frac{1}{2}\ln|u^2+1|+4\arctan u=-\ln|t|+\ln|C|\\
\sqrt{u^2+1}\cdot e^{4\arctan u}=\frac{C}{t}"

Then general solution is

"\sqrt{(\frac{y+2}{x-1})^2+1}\cdot e^{4\arctan \frac{y+2}{x-1}}=\frac{C}{x-1}"